# InvisibleInk: High-Utility and Low-Cost Text Generation with Differential Privacy

### v1

@@ C.1 Adapting InvisibleInk for other Adjacency Notions
The notion of adjacency used is central to the interpretation of the obtained privacy guarantees, as well as to the sensitivity analysis. We use replace-by-null adjacency to derive the privacy guarantees for InvisibleInk in Property 1 and Theorem 2; refer to §B for a full discussion of other adjacency notions and on how they have been used in previous work.
We also state the sensitivity of $f_{\mathsf{DClip}}$ under zero-out adjacency in Property 12.
@@
Let the dataset set $B=|{\bm{R}}|$ be fixed. Then the sensitivity of the map $f_{\mathsf{DClip}}:({\bm{r}}_{1},\ldots,{\bm{r}}_{B})\mapsto\frac{1}{B}\sum_{i=1}^{B}\mathsf{DClip}_{C}({\bm{\phi}}_{i},{\bm{\phi}}_{\mathsf{pub}})$ defined in Property 1, under zero-out adjacency, is
@@
For zero-out adjacency, consider adjacent datasets ${\bm{R}}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{B}\}$ and ${\bm{R}}^{\prime}=\{{\bm{r}}_{1}^{\prime},\ldots,{\bm{r}}_{B}^{\prime}\}$ that differ in their first element where ${\bm{r}}_{1}^{\prime}=\perp$ is the empty string. We need to bound the sensitivity of the map $f_{\mathsf{DClip}}$ from Eq. 9. As in the proof of Property 1, we use shorthand ${\bm{\phi}}_{\mathsf{pub}}:=\phi(\,\cdot\,|\,{\bm{q}},{\bm{x}}_{<t})$ and ${\bm{\phi}}_{i}:=\phi(\,\cdot\,|\,{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})$ are public and private logits respectively for ${\bm{R}}$ and ${\bm{\phi}}_{i}^{\prime}:=\phi(\,\cdot\,|\,{\bm{q}},{\bm{r}}_{i}^{\prime},{\bm{x}}_{<t})$ are private logits for ${\bm{R}}^{\prime}$. Noting that (a) ${\bm{\phi}}_{i}={\bm{\phi}}_{i}^{\prime}$ for $i\neq 1$ by assumption, and (b) ${\bm{\phi}}_{1}^{\prime}=\bm{0}$ under the zero-out adjacency, we upper bound the sensitivity as
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\frac{1}{B}\left\lVert\mathsf{clip}_{C}\big{(}{\bm{\phi}}_{1}-{\bm{\phi}}_{\mathsf{pub}}\big{)}-\mathsf{clip}_{C}\big{(}{\bm{\phi}}_{1}^{\prime}-{\bm{\phi}}_{\mathsf{pub}}\big{)}\right\rVert_{\infty}$
@@
The 2×2$\times$2 × increase in the sensitivity of $f_{\mathsf{DClip}}$ under zero-out adjacency causes a 4×4$\times$4 × increase in the $\rho$-zCDP guarantee of Algorithm 1 for the same parameters and settings as under the replace-by-null adjacency. This highlights the need for careful definition of adjacency notions in DP.
@@ D.1 Task: Synthetic Text Generation
Our goal is to generate high-quality, long-form synthetic data from a given private dataset of references, while under rigorous DP guarantees on the privacy of the references. We model the references ${\bm{R}}=({\bm{r}}_{1},\dots,{\bm{r}}_{|B|})$ as part of the context for next-token generation in the LLM. The query prompts ${\bm{q}}$, contain instructions for text generation including general details about the sensitive dataset such as domain, type of text document, etc. The public and private (here, private prompts mean prompts with sensitive references embedded in-context) query prompts used for the generation of synthetic text samples for various datasets are listed in Tab. 5. Examples of the references for both Yelp and TAB datasets are given in Tab. 6.
Algorithm 1 describes the generation of $1$ text sample from a batch of sensitive references ${\bm{R}}$ of size $B$. To generate multiple synthetic text samples, we partition the entirety of the sensitive reference dataset into $B$-sized chunks. Since the dataset partitions are defined in a completely data-independent manner, we may compose the DP guarantee parallely across batches; see Corollary 11. The theoretical privacy guarantee for the generation of multiple text samples is thus the same as that of $1$ text sample, ie, the overall privacy guarantees are independent of the number of synthetic samples generated and only depend on the maximum number of BPE encoded tokens in each input.
@@ D.2 Datasets
We use both publicly available open-access datasets and licensed datasets: MIMIC IV Clinical Notes dataset [70], Yelp Reviews dataset [72], and the Text Anonymization Benchmark (TAB) [71] dataset. The respective license and data use agreements of the various assets we use have been duly followed throughout this work.
MIMIC-IV Clinical Notes. This dataset is a large, de-identified collection of medical text associated with ICU patients, including discharge summaries, radiology reports, and nursing notes. It is widely used for clinical NLP tasks such as diagnosis classification, named entity recognition, and patient cohort identification. The dataset poses significant privacy challenges due to the presence of sensitive patient information, making it an ideal choice for evaluating privacy-preserving language models. Further, the dataset is not available for open-access and cannot legally be used for pre-training general-purpose LLMs such as those we use in our experiments. This dataset is released under the PhysioNet Credentialed Health Data License 1.5.0, and may be accessed via the PhysioNet website (https://physionet.org/content/mimiciv/3.1/).
For our synthetic text generation task, we use the discharge summaries released as part of the dataset. In particular, we extract the discharge instructions from within these files. These are selected as they often contain succinct summaries of patient symptoms, diagnosis, treatment, as well as aftercare instructions. Further, almost all discharge summaries contain a section on “Discharge Instruction” allowing us to extract almost $311$K valid private references from the dataset. In accordance with the License and Data Use Agreement of the dataset, we do not release the original dataset, processed text samples, or synthetically generated private samples.
Yelp Reviews. We use the processed Yelp Open dataset [72] from [39] released by [36]. This dataset contains $1.9$M of user-generated reviews and ratings for businesses (see Tab. 1 for a precise size), labelled by business category ($10$ categories) and review scores ($5$ possible scores) to give a total of $50$ sub-categories. Although not overtly sensitive, it contains personal opinions and location references and has previously been used to design private synthetic text generation algorithms [39, 36].
@@ D.4 Evaluation Metrics
  • Number of generations and references: Varying between 100 and 1000, as described in Tab. 1. We use an equal number of samples from both the generations and references.
  • Embedding model: Sentence Transformer model [85], version all-mpnet-base-v2.121212 Available at https://huggingface.co/sentence-transformers/all-mpnet-base-v2.
@@
We use the DeBERTa-Med-NER-2 [77] model to label all medical named entities in the generated text. The model recognizes $41$ categories of named-entities of which the text subset of the MIMIC dataset we extract, as described in §D.2, contains only $25$ classes. We further select and count the presence of the most relevant medical named entities; $13$ classes like therapeutic procedure and disease/disorder are included, but domain-agnostic classes like date and time are excluded. The exact details are presented in Tab. 7.
@@
A low perplexity value indicates that the model assigns a higher likelihood to the observed sequence. We wish the generated synthetic text to match the original text dataset in all statistical properties, including perplexity under any given language model. Thus, achieving a small difference in perplexities between the generated and original text is desirable. This is measured by the $\Delta$PPL metric [15], defined as
$\Delta\text{PPL}({\bm{x}})=\left|\text{PPL}({\bm{x}})-\frac{1}{|{\bm{R}}|}\sum_{{\bm{r}}\in{\bm{R}}}\text{PPL}({\bm{r}})\right|\,,$
where ${\bm{R}}$ is the original text dataset comprising sensitive references. We report the average $\Delta\text{PPL}({\bm{x}})$ over all generations of ${\bm{x}}$ (and its confidence intervals, as described §D.6).
@@ D.5 Hyperparameters
InvisibleInk. InvisibleInk, described in Algorithm 1, has the following hyperparameters: $C$ clipping norm, $B$ batch size, $\tau$ sampling temperature and $k$ top-$k$ sampling parameter. For a given $\varepsilon$, the heuristics for choosing optimal hyperparameters are given in §5. To arrive at this heuristic, we conducted a thorough hyperparameter search for a full vocabulary $k=|V|$ variant of the model; the full range of hyperparameters explored for each dataset and model we use is given in Tab. 8. We report results by tuning the sampling temperature for various values of $\varepsilon$ and $B$.
@@
Based on these results (presented in detail in §E) and on the exploration of non-private generation in Fig. 2, we report results for InvisibleInk with $k<|V|$ with a fixed sampling temperature $\tau=1.2$. We explore a range of top-$k$ parameters for various settings; see Tab. 9. Our empirical results in §5 show that setting $k=50$ shows superior performance to all existing baselines, as well as the temperature-tuned full-vocabulary ($k=|V|$) variant of InvisibleInk.
@@
MIMIC TinyLLaMA 1B $B$ $[4,8,16]$
$k$ $[10,50,100]$
LLaMA3.2 1B $B$ $[4,8,16]$
$k$ $[10,50,100]$
LLaMA3 8B $B$ $8$
$k$ $[10,50,100]$
@@
Table 9: Hyperparameters for InvisibleInk ($k<|V|$). We fix $\tau=1.2$ throughout.
Amin et al. [12]. The algorithm proposed by [12] has similar hyperparameters: $C$ clipping norm, $B$ batch size, $\tau_{prv}$ private sampling temperature, $\tau_{pub}$ public sampling temperature, $\theta$ threshold for Sparse Vector Technique (SVT), $\sigma$ noise parameter for SVT. Further, due to the data-dependent privacy accounting proposed in [12, Thm 1], the privacy guarantees depend on the number of times a token is sampled from the aggregated and clipped private logits.
@@
As mentioned in §5, preliminary experimentation showed that the SVT step in [12]’s method selected private tokens $\approx 25\%$ of the time. We select a value of $T_{prv}=$100 for all datasets (where $T_{prv}$ is the maximum number of private tokens allowed) and maintain the total tokens generated at the same levels as the settings for InvisibleInk described in Tab. 1. As seen in the plots in Fig. 4, this does not affect the average generation length of synthetic samples generated by this method. As seen in §5, for a number of settings, this method fails to yield any synthetic text (INF or infeasible in Tab. 2). All results available are reported tuned on the sampling temperature $\tau$.
@@
We use a custom implementation of the method. We note that AdaPMixED has significant (non-GPU) computational overhead compared to other methods due to repeated calculation of optimal mixing parameters; see [18, Alg 1, Line 11]. For generations, we use the same values of $T$ (maximum allowed tokens per generation) as discussed previously; see Tab. 1. We fix some hyperparameters across settings: sampling temperature at $\tau=1$, noise multiplier at $\sigma=0.01$ and mixing parameter for noisy screening $\lambda=10^{-4}$; for all other hyperparameters, we use the recommended default values for the PubMed dataset in [18, Tab 4].

### v2

@@ C.1 Adapting InvisibleInk for other Adjacency Notions
The notion of adjacency used is central to the interpretation of the obtained privacy guarantees, as well as to the sensitivity analysis. We use replace-by-null adjacency to derive the privacy guarantees for InvisibleInk in Property˜1 and Theorem˜2; refer to §B for a full discussion of other adjacency notions and on how they have been used in previous work.
We also state the sensitivity of $f_{\mathsf{DClip}}$ under zero-out adjacency in Property˜12.
@@
Let the dataset set $B=|{\bm{R}}|$ be fixed. Then the sensitivity of the map $f_{\mathsf{DClip}}:({\bm{r}}_{1},\ldots,{\bm{r}}_{B})\mapsto\frac{1}{B}\sum_{i=1}^{B}\mathsf{DClip}_{C}({\bm{\phi}}_{i},{\bm{\phi}}_{\mathsf{pub}})$ defined in Property˜1, under zero-out adjacency, is
@@
For zero-out adjacency, consider adjacent datasets ${\bm{R}}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{B}\}$ and ${\bm{R}}^{\prime}=\{{\bm{r}}_{1}^{\prime},\ldots,{\bm{r}}_{B}^{\prime}\}$ that differ in their first element where ${\bm{r}}_{1}^{\prime}=\perp$ is the empty string. We need to bound the sensitivity of the map $f_{\mathsf{DClip}}$ from Eq.˜10. As in the proof of Property˜1, we use shorthand ${\bm{\phi}}_{\mathsf{pub}}:=\phi(\,\cdot\,|\,{\bm{q}},{\bm{x}}_{<t})$ and ${\bm{\phi}}_{i}:=\phi(\,\cdot\,|\,{\bm{q}},{\bm{r}}_{i},{\bm{x}}_{<t})$ are public and private logits respectively for $\bm{R}$} and ${\bm{\phi}}_{i}^{\prime}:=\phi(\,\cdot\,|\,{\bm{q}},{\bm{r}}_{i}^{\prime},{\bm{x}}_{<t})$ are private logits for $\bm{R}$$^{\prime}$. Noting that (a) ${\bm{\phi}}_{i}={\bm{\phi}}_{i}^{\prime}$ for $i\neq 1$ by assumption, and (b) ${\bm{\phi}}_{1}^{\prime}=\bm{0}$ under the zero-out adjacency, we upper bound the sensitivity as
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\frac{1}{B}\left\lVert\mathsf{clip}_{C}\big({\bm{\phi}}_{1}-{\bm{\phi}}_{\mathsf{pub}}\big)-\mathsf{clip}_{C}\big({\bm{\phi}}_{1}^{\prime}-{\bm{\phi}}_{\mathsf{pub}}\big)\right\rVert_{\infty}$
@@
The $2\times$ increase in the sensitivity of $f_{\mathsf{DClip}}$ under zero-out adjacency causes a $4\times$ increase in the $\rho$-zCDP guarantee of Algorithm˜1 for the same parameters and settings as under the replace-by-null adjacency. This highlights the need for careful definition of adjacency notions in DP.
@@ C.2 Technical Lemmas
C.3 Intuition for Top-$k+$ Sampling
As described in §4, we isolate the contribution of each private logit vector ${\bm{\phi}}_{i}$, following clipping and averaging, as ${\bm{\phi}}^{\mathsf{clip}}_{i}$.
Figure 7: Schematic describing Top-$k+$ sampling. After clipping and aggregating using DClip, the logit values corresponding to each token change by at most $\frac{C}{B}$, i.e., the sensitivity of the private clipped logits. Top-$k+$ sampling, lowers the top-$k$ threshold of the public logits by $\frac{2C}{B}$. The expanded vocabulary $V_{k}^{+}$ captures all tokens in a superset of the top-$k$ vocabularies of the public logits ${\bm{\phi}}_{\mathsf{pub}}$, and all the clipped private logits ${\bm{\phi}}^{\mathsf{clip}}_{i}$.
In Fig.˜7, the logit values of each token in ${\bm{\phi}}^{\mathsf{clip}}_{i}$ differ from those in ${\bm{\phi}}_{\mathsf{pub}}$ by at most $\frac{C}{B}$. Further, the top-$k$ sets of both these logits may differ as token ordering around the top-$k$ threshold changes in the following two ways: (i) tokens in the top-$k$ set of ${\bm{\phi}}_{\mathsf{pub}}$ can drop out of the top-$k$ set of ${\bm{\phi}}^{\mathsf{clip}}_{i}$ as their logit value decreases, and (ii) tokens outside the top-$k$ set of ${\bm{\phi}}_{\mathsf{pub}}$ can enter the top-$k$ set of ${\bm{\phi}}^{\mathsf{clip}}_{i}$ as their logit value increases.
All tokens which exit the top-$k$ set of ${\bm{\phi}}_{\mathsf{pub}}$ due to (i), must still lie within $\frac{C}{B}$ of the top-$k$ threshold of ${\bm{\phi}}_{\mathsf{pub}}$. However, this could potentially decrease the top-$k$ threshold of ${\bm{\phi}}_{\mathsf{pub}}$ by at most $\frac{C}{B}$ when moving to ${\bm{\phi}}^{\mathsf{clip}}_{i}$. Thus, all tokens which enter the top-$k$ set of ${\bm{\phi}}^{\mathsf{clip}}_{i}$ due to (ii) must lie within $\frac{2C}{B}$ of the top-$k$ threshold of ${\bm{\phi}}_{\mathsf{pub}}$. Thus, lowering the top-$k$ threshold of ${\bm{\phi}}_{\mathsf{pub}}$ by $\frac{2C}{B}$ captures all tokens of interest, as shown by the toy example in Fig.˜7.
@@ D.1 Task: Synthetic Text Generation
Our goal is to generate high-quality, long-form synthetic data from a given private dataset of references, while under rigorous DP guarantees on the privacy of the references. We model the references ${\bm{R}}=({\bm{r}}_{1},\dots,{\bm{r}}_{|B|})$ as part of the context for next-token generation in the LLM. The query prompts $\bm{q}$}, contain instructions for text generation including general details about the sensitive dataset such as domain, type of text document, etc. The public and private (here, private prompts mean prompts with sensitive references embedded in-context) query prompts used for the generation of synthetic text samples for various datasets are listed in Tab.˜5. Examples of the references for both Yelp and TAB datasets are given in Tab.˜6.
Algorithm˜1 describes the generation of $1$ text sample from a batch of sensitive references $\bm{R}$} of size $B$. To generate multiple synthetic text samples, we partition the entirety of the sensitive reference dataset into $B$-sized chunks. Since the dataset partitions are defined in a completely data-independent manner, we may compose the DP guarantee parallely across batches; see Corollary˜11. The theoretical privacy guarantee for the generation of multiple text samples is thus the same as that of $1$ text sample, i.e., the overall privacy guarantees are independent of the number of synthetic samples generated and only depend on the maximum number of BPE encoded tokens in each input.
@@ D.2 Datasets
We use both publicly available open-access datasets and licensed datasets: MIMIC IV Clinical Notes dataset [78, 79], Yelp Reviews dataset [81], and the Text Anonymization Benchmark (TAB) [80] dataset. The respective license and data use agreements of the various assets we use have been duly followed throughout this work.
MIMIC-IV Clinical Notes. This dataset is a large, de-identified collection of medical text associated with ICU patients, including discharge summaries, radiology reports, and nursing notes. It is widely used for clinical NLP tasks such as diagnosis classification, named entity recognition, and patient cohort identification. The dataset poses significant privacy challenges due to the presence of sensitive patient information, making it an ideal choice for evaluating privacy-preserving language models. Further, the dataset is not available for open-access and cannot legally be used for pre-training general-purpose LLMs such as those we use in our experiments. This dataset is released under the PhysioNet Credentialed Health Data License 1.5.0, and may be accessed via the PhysioNet website (https://physionet.org/content/mimic-iv-note/2.2/).
For our synthetic text generation task, we use the discharge summaries released as part of the dataset. In particular, we extract the discharge instructions from within these files. These are selected as they often contain succinct summaries of patient symptoms, diagnosis, treatment, as well as aftercare instructions. Further, almost all discharge summaries contain a section on “Discharge Instruction” allowing us to extract almost 311311K valid private references from the dataset. In accordance with the License and Data Use Agreement of the dataset, we do not release the original dataset, processed text samples, or synthetically generated private samples.
Yelp Reviews. We use the processed Yelp Open dataset [81] from [44] released by [41]. This dataset contains 1.91.9M of user-generated reviews and ratings for businesses (see Tab.˜1 for a precise size), labelled by business category ($10$ categories) and review scores ($5$ possible scores) to give a total of $50$ sub-categories. Although not overtly sensitive, it contains personal opinions and location references and has previously been used to design private synthetic text generation algorithms [44, 41].
@@
Note on Choice of Dataset. We note that several common datasets used in prior work (such as AGNews, TREC, DBPedia, WikiMovies, etc.) are incompatible with our setting of long-form text generation due to their relatively shorter texts ($\leq 200$ tokens). Further, several of these datasets (like IMDb, etc.) are well-represented in the pretraining data of most LLMs. Since publicly prompted LLMs can generate high-quality synthetic text for such datasets, a high degree of overlap between the pretraining and target data distributions skews the conclusions of private data analysis [82]. Thus, we prefer using benchmarks that are sufficiently out-of-distribution relative to the pretraining data and preferably contain real privacy-sensitive data.
Accordingly, we use the MIMIC IV Notes dataset as the primary dataset for experiments, since it contains real patient discharge notes with sensitive medical data, and cannot legally be used for pretraining LLMs as per its data use agreement. Similarly, the TAB dataset contains sensitive identifying information (demographic traits, spatio-temporal markers, biographical details, etc.) from European Court of Human Rights cases. We use the Yelp dataset as a representative dataset that overlaps significantly with the pretraining data (similar to the other datasets discussed above).
@@ D.4 Evaluation Metrics
  • Number of generations and references: Varying between 100 and 1000, as described in Tab.˜1. We use an equal number of samples from both the generations and references.
  • Embedding model: Sentence Transformer model [95], version all-mpnet-base-v2.111111Available at https://huggingface.co/sentence-transformers/all-mpnet-base-v2.
@@
We use the DeBERTa-Med-NER-2 [88] model to label all medical named entities in the generated text. The model recognizes $41$ categories of named-entities of which the text subset of the MIMIC dataset we extract, as described in §D.2, contains only $25$ classes. We further select and count the presence of the most relevant medical named entities; $13$ classes like therapeutic procedure and disease/disorder are included, but domain-agnostic classes like date and time are excluded. The exact details are presented in Tab.˜7.
@@
[…]
$\Delta\text{PPL}=\left|\frac{1}{|{\bm{X}}|}\sum_{{\bm{x}}\in{\bm{X}}}\text{PPL}({\bm{x}})-\frac{1}{|{\bm{R}}|}\sum_{{\bm{r}}\in{\bm{R}}}\text{PPL}({\bm{r}})\right|\,$,
where $\bm{R}$} is the original dataset of sensitive references and $\bm{X}$} is the synthetically generated dataset.
@@ D.5 Hyperparameter Selection
InvisibleInk. InvisibleInk, described in Algorithm˜1, has the following hyperparameters: the clipping norm $C$, the batch size $B$, the sampling temperature $\tau$ and the top-$k$ sampling parameter. For a given $\varepsilon$, the heuristics for choosing optimal hyperparameters are given in §5. To arrive at this heuristic, we conducted a thorough hyperparameter search for a full vocabulary $k=|V|$ variant of the model; the full range of hyperparameters explored for each dataset and model we use is given in Tab.˜8. We report results by tuning the sampling temperature for various values of $\varepsilon$ and $B$.
@@
Based on these results (presented in detail in §E) and on the exploration of non-private generation in Fig.˜2, we report results for InvisibleInk with a fixed sampling temperature $\tau=1.1$ for $k<|V|$ and $\tau=1.0$ for $k=|V|$. We explore a range of top-$k$ parameters for various settings; see Tab.˜9. Our empirical results in §5 show that setting k=50100k=50-100 shows superior performance to all existing baselines, as well as the full-vocabulary ($k=|V|$) variant of InvisibleInk.
@@
MIMIC TinyLLaMA 1B $B$ $[4,8,16,32]$
$k$ $[10,50,100,500]$
$\tau$ $[1.0,1.1.,1.2]$
LLaMA3.2 1B $B$ 88
$k$ $[10,50,100,500]$
$\tau$ $1.1$
LLaMA3 8B $B$ 88
$k$ $[10,50,100,500,1000,5000]$
@@
Table 9: Hyperparameters for InvisibleInk ($k<|V|$). We set $\tau=1.2$ wherever unspecified.
Amin et al. [13]. The algorithm proposed by [13] has similar hyperparameters: $C$ clipping norm, $B$ batch size, $\tau_{prv}$ private sampling temperature, $\tau_{pub}$ public sampling temperature, $\theta$ threshold for Sparse Vector Technique (SVT), $\sigma$ noise parameter for SVT. Further, due to the data-dependent privacy accounting proposed in [13, Thm 1], the privacy guarantees depend on the number of times a token is sampled from the aggregated and clipped private logits.
@@
As mentioned in §5, preliminary experimentation showed that the SVT step in [13]’s method selected private tokens $\approx 25\%$ of the time. We select a value of $T_{prv}=$100 for all datasets (where $T_{prv}$ is the maximum number of private tokens allowed) and maintain the total tokens generated at the same levels as the settings for InvisibleInk described in Tab.˜1. As seen in the plots in Fig.˜4, this does not affect the average generation length of synthetic samples generated by this method. As seen in §5, for a number of settings, this method fails to yield any synthetic text (INF or infeasible in Tab.˜2). All results available are reported tuned on the sampling temperature $\tau$.
@@
We use a custom implementation of the method, modifying it to use the replace-by-null adjacency; see §B.3 & §E.5. We note that AdaPMixED has significant (non-GPU) computational overhead compared to other methods due to repeated calculation of optimal mixing parameters; see [19, Alg 1, Line 11]. We implement this using the brentq optimizer from the scipy Python package. For generations, we use the same values of $T$ (maximum allowed tokens per generation) as discussed previously; see Tab.˜1. We fix some hyperparameters across settings: sampling temperature at $\tau=1$, noise multiplier at $\sigma=0.01$ and mixing parameter for noisy screening $\lambda=10^{-4}$; for all other hyperparameters, we use the recommended default values for the PubMed dataset in [19, Tab 4].